Determine the mass of Mg3N2 that contain 3.97 x 1021 N atoms.
Answer #1

solution: to solve the given problem, we have to use these two formulas:

i) Number of moles = Number of atoms in Mg3N2 / 6.022 × 1023

ii) Number of moles = Mass (in grams) / Molar mass of Mg3N2.

Given: No of atoms of Mg3N2 = 3.97 × 1021 atoms.

------------------

step 1):- First we have to calculate the number of moles of Mg3N2 using formula i), as given below:-

Number Of moles = 3.97 × 1021/ 6.022 × 1023 = 0.66 × 10-2 mol

​​​​​​step 2):- Now, we have to calculate molar mass of Mg3N2 = (3 × Atomic mass of Mg) + (2 × Atomic mass of N) = (3 × 24.305) + (2 × 14.0067) = 100.94 g/mol

Step 3):- Now, putting values of number of moles and molar mass in formula ii) ,we get ,

0.66 × 10​​​​​-2 = Mass of Mg3N2 / 100.94

Mass of Mg3N2 = 0.66 × 10-2 × 100.94 = 0.66 grams

Hence, mass of Mg3N2 is = 0.66 g

​​​​​​